# Jigsaw-R1: A Study of Rule-based Visual Reinforcement Learning with Jigsaw Puzzles

### v2

@@ 5 Experiments
Visual reasoning models. Recent advancements from OpenAI, particularly the o3 and o4-mini models (OpenAI, 2025b), have shown significant promise in reasoning with images for enhanced perception. While our work does not incorporate these visual reasoning models (Li et al., 2025a; Liu et al., 2025f; Qi et al., 2025; Su et al., 2025; Wang et al., 2025c), we believe jigsaw puzzles are an ideal candidate for exploring rule-based visual RL in this context due to their inherent reliance on image-based reasoning. For an early exploration, we conduct small-scale experiments using the ChatGPT console, where these models are equipped with tool-use capabilities. Our preliminary experiments indicate that OpenAI o3 can effectively solve 2x2 jigsaw puzzles, substantially outperforming other models considered in this paper. However, it still faces challenges with more complex puzzles (e.g. 3x3), highlights areas for further investigation.
@@
Multimodal generative models. Our study does not consider models capable of both understanding and generating multimodal content (Chen et al., 2025c; e; Hurst et al., 2024; Team, 2024; Wu et al., 2024; Xie et al., 2025a). A promising future research direction involves integrating our approach with these advanced models. Enabling them to generate their own inputs could reduce dependence on external datasets like COCO and create an autonomous environment to learn from experience (Silver & Sutton, 2025).
@@
Test-time training. We have demonstrated that aligning the training dataset with the target domain can yield enhanced performance. Jigsaw puzzles inherently provide readily available annotations, making them suitable for direct training on the test set during test time. Therefore, exploring the use of jigsaw puzzles as a technique for test-time training (Akyurek et al., 2024; Behrouz et al., 2024; Zhu et al., 2024; Zuo et al., 2025) presents an interesting avenue for future work.
@@
Other pretext tasks. While this work primarily focuses on jigsaw puzzles as the pretext task, numerous alternatives exist and worth exploration (Gidaris et al., 2018; Gui et al., 2024). In particular, future research could extend our approach to pretext tasks in other modalities, including text (Lan et al., 2020), video (Ahsan et al., 2019; Kim et al., 2019; Wang et al., 2020), audio (Carr et al., 2021), point clouds (Poursaeed et al., 2020), and tabular data (Lee et al., 2024).
@@
Other RL algorithms. We exclusively employ GRPO in our current experiments, leaving other promising algorithms such as PPO (Schulman et al., 2017), DPO (Rafailov et al., 2023) and Reinforce++ (Hu, 2025) unexplored. Furthermore, investigating recent advancements and variations of GRPO, including DAPO (Yu et al., 2025b), Dr. GRPO (Liu et al., 2025d), GPG (Chu et al., 2025), and NoisyRollout (Liu et al., 2025b), could offer valuable insights and potential performance gains.
@@ Acknowledgements
Appendix A Full Jigsaw Puzzles
@@
Appendix B Box Jigsaw Puzzles
This task evaluates a MLLM’s ability to reason about spatial relationships between image patches and to visually ground its understanding by identifying the correct bounding box.
The process begins by dividing an input image into an $m{\scriptstyle\times}n$ grid of regions. Each region is assigned a unique position index. From each of these $mn$ regions, a patch of equal size is randomly selected, resulting in $mn$ equally-sized patches. Next, a target position index, $i$ (where $1\leq i\leq mn$), is randomly chosen. All $mn$ patches are then shuffled. A key constraint during shuffling is that the patch originally from position $i$ must be relocated to a different position. The MLLM’s task is to provide the bounding box coordinates for the patch that originally belonged in position $i$. The accuracy reward is measured by the Intersection over Union (IoU) between the predicted and ground truth bounding boxes.
As shown in Table 8, training MLLMs on this task can promote generalization to ScreenSpot Cheng et al. (2024), a downstream task requiring visual grounding within a Graphical User Interface (GUI) environment.
@@ Appendix B Box Jigsaw Puzzles
Appendix C Qwen2-VL-2B-Base
@@
The training dynamics of Qwen2-VL-2B-Base on 2x1 pair jigsaw puzzles are illustrated in Figure 4. While rewards show a rapid initial increase, mirroring trends seen in instruction-tuned models, the improvement in accuracy reward is largely superficial. It primarily reflects the model learning to adhere to the specified output format, which allows for the extraction of a final answer, rather than indicating a genuine enhancement in its capabilities to solve jigsaw puzzles. Subsequently, the accuracy reward stagnates, not surpassing 0.5, which is equivalent to the performance of random guessing. This leads us to hypothesize that jigsaw puzzles present a significant challenge for the base model, potentially requiring a substantially extended training period to achieve meaningful performance gains.
@@
Similar to observations with instruction-tuned models, the keywords are already present in the model before training begins (at step 0). This aligns with findings in textual domain (Liu et al., 2025d). Besides, while these keywords appear more often as training progresses, their overall occurrence remains relatively infrequent.
@@
Appendix D Keywords
@@
Appendix E Prompts
This section provides prompts for different variants of jigsaw puzzles, using a 2x2 puzzle as an example.
@@ Appendix E Prompts
Appendix F Qwen2.5-VL-3B Examples

### v3

@@ 5 Experiments
Skill alignment. The efficacy of jigsaw puzzles is contingent upon skill alignment. The primary skill cultivated is spatial reasoning, which is readily transferable to applications where the spatial relationship between objects is critical, such as the downstream tasks in our experiments. In contrast, for tasks less reliant on this visual ability (e.g. mathematical reasoning), the training may yield diminishing returns. Therefore, we recommend integrating jigsaw puzzles with other training methods for broader gains rather than using them in isolation. Determining whether to always include jigsaw puzzles and similar pretext tasks (e.g. image rotation, as discussed in Appendix˜C) in training data requires careful, data-centric ablation studies (Li et al., 2025d), which we leave for future work. Encouragingly, recent studies have already shown that incorporating jigsaw puzzles into training can enhance visual reasoning capabilities of MLLMs (Li et al., 2025a; Wu et al., 2025; Zeng et al., 2025).
Data contamination. It is important to consider our findings in the context of potential data contamination, a well-known challenge in the evaluation of MLLMs. Because the models are pre-trained on vast and opaque web-scale datasets, their training data likely overlaps with our evaluation benchmarks. Consequently, the development of evaluation protocols that are robust against such contamination represents a critical and unresolved challenge for the research community.
@@
Visual reasoning models. Recent advancements from OpenAI, particularly the o3 and o4-mini models (OpenAI, 2025b), have shown significant promise in reasoning with images for enhanced perception. While our work does not incorporate these visual reasoning models (Li et al., 2025b; Liu et al., 2025g; Qi et al., 2025; Su et al., 2025; Wang et al., 2025c), we believe jigsaw puzzles are an ideal candidate for exploring rule-based visual RL in this context due to their inherent reliance on image-based reasoning. For an early exploration, we conduct small-scale experiments using the ChatGPT console, where these models are equipped with tool-use capabilities. Our preliminary experiments indicate that OpenAI o3 can effectively solve 2x2 jigsaw puzzles, substantially outperforming other models considered in this paper. However, it still faces challenges with more complex puzzles (e.g. 3x3), highlights areas for further investigation.
@@
Multimodal generative models. Our study does not consider models capable of both understanding and generating multimodal content (Chen et al., 2025c; e; Hurst et al., 2024; Team, 2024; Wu et al., 2024; Xie et al., 2025a). A promising future research direction involves integrating jigsaw puzzles with these advanced models and let them to generate their own inputs. This could reduce dependence on external datasets like COCO and create an autonomous environment to learn from experience (Silver & Sutton, 2025).
@@
Test-time training. Jigsaw puzzles inherently provide readily available annotations, making them suitable for direct training on the test set during test time. We have demonstrated that aligning the training dataset with the target domain can yield enhanced performance. Therefore, exploring the use of jigsaw puzzles as a technique for test-time training (Akyurek et al., 2024; Behrouz et al., 2024; Zhu et al., 2024; Zuo et al., 2025) presents an interesting avenue for future work.
@@
Other pretext tasks. While this work primarily focuses on jigsaw puzzles as the pretext task, numerous alternatives exist and worth exploration (Gidaris et al., 2018; Gui et al., 2024). As a preliminary step in this direction, we include a study on image rotation in Appendix˜C. Besides, future research could extend our approach to pretext tasks in other modalities, including text (Lan et al., 2020), video (Ahsan et al., 2019; Kim et al., 2019; Wang et al., 2020), audio (Carr et al., 2021), point clouds (Poursaeed et al., 2020), and tabular data (Lee et al., 2024).
@@
Other RL algorithms. We exclusively employ GRPO in our current experiments, leaving other RL algorithms such as PPO (Schulman et al., 2017), DPO (Rafailov et al., 2023) and Reinforce++ (Hu, 2025) unexplored. Furthermore, investigating recent advancements and variations of GRPO, including DAPO (Yu et al., 2025b), Dr. GRPO (Liu et al., 2025e), GPG (Chu et al., 2025), and NoisyRollout (Liu et al., 2025c), could offer valuable insights and potential performance gains.
@@ Acknowledgements
Appendices
Appendix A Other Jigsaw Puzzles
A.1 Full Jigsaw Puzzles
@@
A.2 Box Jigsaw Puzzles
Unlike other jigsaw puzzle tasks (i.e., full and pair), this task evaluates a MLLM’s ability to both reason about spatial relationships between image patches and to visually ground its understanding by identifying the correct bounding box.
The process begins by dividing an input image into an $m{\scriptstyle\times}n$ grid of regions. Each region is assigned a unique position index. From each of these $mn$ regions, a patch of equal size is randomly selected, resulting in $mn$ equally-sized patches. Next, a target position index, $i$ (where $1\leq i\leq mn$), is randomly chosen. All $mn$ patches are then shuffled. A key constraint during shuffling is that the patch originally from position $i$ must be relocated to a different position. The MLLM’s task is to provide the bounding box coordinates for the patch that originally belonged in position $i$. Further details on the prompts used can be found in Appendix˜I.
The accuracy reward is measured by the Intersection over Union (IoU) between the predicted and ground truth bounding boxes. The format reward, consistent with other jigsaw puzzle tasks, requires that the final answer be extractable in the prescribed format.
As shown in Table 8, training MLLMs on this task can promote generalization to ScreenSpot (Cheng et al., 2024), a downstream task requiring visual grounding within a Graphical User Interface (GUI) environment.
@@ A.2 Box Jigsaw Puzzles
Appendix B Other Reward Designs
B.1 Binary Accuracy Reward
For full questions with fine-grained reward signals, we default to using fractional rewards. As an alternative, we test binary rewards, similar to pair questions. To compare these two designs, we evaluate the generalization performance of Qwen2.5-VL-3B (non-thinking) on downstream tasks. As shown in Table˜9, the fractional reward yields slightly better results on larger jigsaw puzzles.
Table 9: Averaged downstream task performance of Qwen2.5-VL-3B (non-thinking) when trained on full questions with different accuracy rewards.
Accuracy Reward 2x1 3x1 4x1 2x2
Binary 72.65 72.70 72.16 71.55
Fractional 72.65 72.77↑0.07 72.77↑0.61 71.94↑0.39
B.2 Format Reward Weight
We assign weights of 1.0 and 0.5 to the accuracy and format rewards, respectively. We also perform an ablation study and explore alternative values such as 0 and 1 for the format reward. The downstream task performance of Qwen2.5-VL-3B (non-thinking) is measured for each configuration. As presented in Table˜10, the choice of the format reward weight has only a minor effect on the final performance.
Table 10: Averaged downstream task performance of Qwen2.5-VL-3B (non-thinking) when trained with different format reward weights.
Weight 0 0.5 1.0
Performance 72.97 73.18 73.06
Appendix C Image Rotation
C.1 Task Design
As an alternative pretext task, we also conduct a preliminary study with image rotation. In this setup, an input image is randomly rotated clockwise by one of a predefined set of angles. Given a size parameter $n$, the possible rotation angles are evenly distributed within a 360 range. We frame this as an $n$-choice question where the model needs to identify the applied rotation angle. For instance, if nn=4, the possible choices are 0, 90, 180, and 270. As with jigsaw puzzles, the difficulty can be adjusted by varying the value of $n$. The specific prompts are available in Appendix˜I.
Consistent with our jigsaw puzzle experiments, we investigate both thinking and non-thinking. Besides, the reward design and training hyperparameters are identical to those used for the jigsaw puzzles. For example, a correct choice receives an accuracy reward of 1 (0 otherwise), and a response from which a letter choice can be extracted receives a format reward of 0.5 (0 otherwise).
C.2 Task Performance
We first train Qwen2.5-VL-3B on the rotation task with 4 choices and then evaluate its performance across a range of sizes. The results are presented in Table˜11. Similar to our findings with jigsaw puzzles, the model initially struggles but masters the task effectively after fine-tuning. Moreover, the model demonstrates the ability to generalize to more complex configurations (i.e., larger rotation sizes) not encountered during training.
Table 11: Evaluation results on image rotation with different sizes. For thinking and non-thinking of the same model, the better result is underlined.
Thinking
Method 4 8 12 15 20 AVG
Random 25.00 12.50 8.33 6.67 5.00 11.50
Qwen2.5-VL-3B 29.80 15.10 $\underline{9.10}$ $\underline{6.50}$ 5.30 13.16
+ Rotation-R1 91.80↑62.00 39.10↑24.00 27.90↑18.80 10.20↑3.70 12.20↑6.90 36.23↑23.07
Non-thinking
Method 4 8 12 15 20 AVG
Random 25.00 12.50 8.33 6.67 5.00 11.50
Qwen2.5-VL-3B $\underline{30.60}$ $\underline{16.80}$ 8.00 6.40 $\underline{5.50}$ $\underline{13.45}$
+ Rotation-R1 $\underline{95.00}$↑64.40 $\underline{48.20}$↑31.40 $\underline{32.60}$↑24.60 $\underline{12.50}$↑6.10 $\underline{18.20}$↑12.70 $\underline{41.30}$↑27.85
C.3 Generalization
We train the model using various rotation sizes and assess its generalization capabilities on downstream tasks, with the results shown in Table˜12. Again, consistent with the jigsaw puzzle experiments, the model trained on image rotation generalizes to unseen tasks, and the degree of generalization varies based on rotation sizes. Furthermore, Qwen2.5-VL-3B learns and generalizes both with and without explicit reasoning, though it tends to benefit from direct answering.
A comparison between the generalization performance of jigsaw puzzles (Table˜3) and rotation tasks (Table˜12) reveals that the former yields more effective generalization. However, we do not claim that jigsaw puzzles are an inherently superior pretext task. Our primary focus on jigsaw puzzles in this paper is to maintain a controlled experimental setup for studying various aspects of rule-based visual RL. Nevertheless, we believe that more effective generalization could be achieved by integrating multiple visual pretext tasks, a strategy analogous to contemporary practices in linguistic domains where various logic puzzles are jointly employed to enhance reasoning capabilities (Chen et al., 2025b; Liu et al., 2025a; b).
Table 12: Averaged downstream task performance of Qwen2.5-VL-3B when trained on different rotation sizes. For thinking and non-thinking of the same model, the better result is underlined.
Thinking
Baseline 4 8 12 15 20
57.64 64.19 62.87 63.58 62.77 63.38
Non-thinking
Baseline 4 8 12 15 20
$\underline{69.59}$ $\underline{71.37}$ $\underline{71.33}$ $\underline{72.55}$ $\underline{72.74}$ $\underline{72.00}$
Appendix D Training Costs
On-policy RL algorithms like GRPO are computationally demanding, as they require generating multiple roll-outs during training. This cost is significantly exacerbated when the model must also produce an explicit reasoning process. We measure these training costs on a cluster of eight 64GB AMD MI250X GPUs. As detailed in Table˜13, RL is substantially more expensive than SFT, especially when the model is required to articulate its reasoning.
Table 13: A comparison of the time (in seconds) required to complete a single training step between RL and SFT for both thinking and non-thinking.
Thinking
Method RL SFT
Time (s) 297.51 9.41
Non-thinking
Method RL SFT
Time (s) 51.57 8.66
Appendix E Training Steps
We compare thinking and non-thinking across training steps. As Table˜14 illustrates, extending the number of training iterations consistently improves downstream generalization for both approaches, albeit at a decreasing rate. This confirms the value of additional training steps when computationally feasible. Notably, thinking consistently outperforms non-thinking across different training steps, aligning with our observations.
Table 14: Averaged downstream task performance of Qwen2.5-VL-3B when trained with different number of steps. For thinking and non-thinking of the same model, the better result is underlined.
Thinking
Training Steps 1,000 2,000
Performance 60.86 61.24
Non-thinking
Training Steps 1,000 2,000
Performance $\underline{72.47}$ $\underline{73.18}$
Appendix F Qwen2-VL-2B-Base
@@
The training dynamics of Qwen2-VL-2B-Base on 2x1 pair jigsaw puzzles are illustrated in Figure˜4. While rewards show a rapid initial increase, mirroring trends seen in instruction-tuned models, the improvement in accuracy reward is largely superficial. It primarily reflects the model learning to adhere to the specified output format, which allows for the extraction of a final answer, rather than indicating a genuine enhancement in its capabilities to solve jigsaw puzzles. Subsequently, the accuracy reward stagnates, not surpassing 0.5, which is equivalent to the performance of random guessing. This leads us to hypothesize that jigsaw puzzles present a significant challenge for the base model, potentially requiring a substantially extended training period to achieve meaningful performance gains.
@@
Similar to observations with instruction-tuned models, the keywords are already present in the model before training begins (at step 0). This aligns with findings in textual domain (Liu et al., 2025e). Besides, while these keywords appear more often as training progresses, their overall occurrence remains relatively infrequent.
@@
Appendix G Few-shot Learning
This section details the few-shot learning performance of Qwen2.5-VL-3B (non-thinking) on 2x1 pair jigsaw puzzles. For each test run, we randomly sample a fixed set of few-shot examples from the training set. The process is repeated five times with different random seeds to compute the mean and standard deviation.
As shown in Table˜15, the model’s performance slightly increases with two shots but then declines as more examples are added. The initial improvement likely serves to provide additional contextual clues, whereas the subsequent drop suggests the model struggles with longer contexts. Ultimately, the performance ceiling of approximately 53.7% highlights that jigsaw puzzles remain a significant challenge for the model.
Table 15: Evaluation results on 2x1 pair jigsaw puzzles of Qwen2.5-VL-3B (non-thinking) with different number of shots.
Shot 0 2 4 8 16
Performance 52.20 $53.74\pm 1.08$ $50.56\pm 0.86$ $49.50\pm 0.41$ $50.30\pm 1.03$
Appendix H Reasoning Patterns
H.1 Keywords
@@
H.2 Content Analysis
Except keyword matching, we also use GPT-4.1 to analyze whether a reasoning process exhibits the four cognitive behaviors, adopting prompts from Gandhi et al. (2025). As shown in Figure˜5, the frequencies of verification and backtracking increase throughout training, mirroring the observation from keyword matching. In contrast, the frequencies of subgoal setting and backward chaining remain largely unchanged.
Figure 5: Evolution of cognitive behavior frequencies throughout the training process.
Appendix I Prompts
I.1 Full Jigsaw Puzzles
@@ I.3 Box Jigsaw Puzzles
I.4 Image Rotation
Image Rotation with 4 choices (thinking) The input image has been rotated clockwise by an angle that is a multiple of 90. Your task is to determine the angle of clockwise rotation. Select the correct answer from the following 4 choices: (A) 0 (B) 180 (C) 90 (D) 270 First, output the thinking process within <think> </think> tags. Then, provide the final answer within <answer> </answer> tags. The final answer should be a single letter.
Image Rotation with 4 choices (non-thinking) The input image has been rotated clockwise by an angle that is a multiple of 90. Your task is to determine the angle of clockwise rotation. Select the correct answer from the following 4 choices: (A) 0 (B) 180 (C) 90 (D) 270 Directly output the final answer. The final answer should be a single letter.
Appendix J Examples
J.1 Qwen2.5-VL-3B